Consider the system obtained from the augmented matrix below abde|cf choose the correct statements 1 the system has no solution if ae=bd


1
Expert's answer
2021-05-07T12:30:37-0400

ae=bd is a unique solution of the given system as (1,1) is the ... Consider a system of linear equations 2x + 5y = 7 and 3x + 2y = 6 ... have been made on the augmented matrix,

Not only that but also, Systems of Linear Equations. Augmented Matrix Notation. Another interesting rewriting is x. (1. 2. ) + y. ( 1. −1. ) = (27. 0. ) . This tells us that ..that the system has no solution if ae=bd